If a firms debt-to-equity ratio is 3.8 , what is its total debt to total asset ratio? Show your answer to the nearest .1%. If your answer is 45.6%, enter your answer as 45.6 rather than 45.6% or .456.
Answer #1

what is its total debt to total asset ratio

=debt/(equity+debt)

=3.8/(1+3.8)

=79.2%

the above is answer..
Answer #2

The total debt to total asset ratio is calculated by dividing the total debt of the firm by its total assets. Given that the debt-to-equity ratio is 3.8, we can calculate the total debt to total asset ratio as follows:

Debt-to-equity ratio = Total debt / Total equity

Since the debt-to-equity ratio is given as 3.8, we can write:

3.8 = Total debt / Total equity

Now, let's rearrange the equation to solve for Total debt:

Total debt = 3.8 * Total equity

To calculate the total debt to total asset ratio, we need to express the total debt as a proportion of total assets. We can do this by dividing the total debt by the sum of total debt and total equity:

Total debt to total asset ratio = Total debt / (Total debt + Total equity)

Substituting the value of Total debt from the previous equation, we get:

Total debt to total asset ratio = (3.8 * Total equity) / (3.8 * Total equity + Total equity) = (3.8 * Total equity) / (4.8 * Total equity) = 3.8 / 4.8

Calculating this ratio, we find:

Total debt to total asset ratio = 0.791666...

Rounding this value to the nearest 0.1%, the total debt to total asset ratio is approximately 79.2%.

Therefore, the total debt to total asset ratio is 79.2%.

answered by: Hydra Master